# P2 Law: Scaling Law for Post-Training After Model Pruning

### v2

@@ 5.2 Scaling Law
We are the first to investigate the scaling law for the post-training after model pruning, and we propose the P2 Law as a scaling law for this process. Compared to the OpenAI scaling law, the Chinchilla scaling law demonstrates superior performance (detailed in Appendix D). Therefore, we adopt the Chinchilla scaling law as the foundational parameterization for our P2 Law.
@@ 6 Conclusion
The P2 Law presented in this paper, derived from experiments on the Llama-3 and Qwen-2.5 series models, may not be directly applicable to other architectures, such as the currently popular Mixture of Experts (MoE) architecture. Therefore, we plan to conduct more experiments across a broader range of model architectures to enhance the generalizability of our findings.
@@ B.2 Width Pruning
In the post-training process, We train this 2:4 sparse model pruned by SparseGPT. Inspired by SP-LoRA Anonymous (2024), during the post-training process, we combine the updated weight $\Delta\tilde{W}^{t}$ from each training iterate $t$ with the mask $M$ to obtain the weight after update $\tilde{W}^{t}$, ensuring the model’s sparsity:
@@
Appendix C Additional Actual Loss Curves
The additional post-training loss curves for models pruned by width pruning or for the Qwen-2.5 series models are provided in Figures 1011, 12 and 14.
Appendix D Comparison with OpenAI Scaling Law
@@
Appendix E Parameter Values of Fitted Parameterizations
@@
Appendix F Additional Loss Curves Derived by P2 Law
The additional loss curves derived by P2 Law are shown in the Figure 151617 and 18.
Appendix G Patterns of the Llama-3 Series Models in Terms of Width
@@
Appendix H Additional Generalization Loss Curves
We present the additional dataset size generalization predicted loss curves in the Figure 20212223 and 24, model size generalization predicted loss curves in the Figure 25 and pruning rate generalization predicted loss curves in the Figure 26 and 27.
@@ Appendix H Additional Generalization Loss Curves
Llama-3 series $\mathcal{L}_{1}$ 0.02 5.94 0.14 -1.57 0.23 1.08 0.29 0.05 5.86 -2.52 -1.68 0.08 -0.97 0.38 38.26 0.87 2.49 26.53 0.37 0.05
@@
(a) Loss curves derived by P2 Law for validating generalization on dataset size (Llama-3.2-1B pruned by depth pruning).
(b) Loss curves derived by P2 Law for validating generalization on dataset size (Llama-3.2-3B pruned by depth pruning).
(c) Loss curves derived by P2 Law for validating generalization on dataset size (Llama-3.1-8B pruned by depth pruning).
@@
Figure 20: Loss curves derived by P2 Law for validating generalization on dataset size (Llama-3 series models pruned by depth pruning).
(a) Loss curves derived by P2 Law for validating generalization on dataset size (Llama-3.2-1B pruned by width pruning).
(b) Loss curves derived by P2 Law for validating generalization on dataset size (Llama-3.2-3B pruned by width pruning).
@@
Figure 21: Loss curves derived by P2 Law for validating generalization on dataset size (Llama-3 series models pruned by width pruning).
(a) Loss curves derived by P2 Law for validating generalization on dataset size (Qwen-2.5-0.5B pruned by depth pruning).
(b) Loss curves derived by P2 Law for validating generalization on dataset size (Qwen-2.5-1.5B pruned by depth pruning).
(c) Loss curves derived by P2 Law for validating generalization on dataset size (Qwen-2.5-3B pruned by depth pruning).
@@
Figure 22: Loss curves derived by P2 Law for validating generalization on dataset size (Qwen-2.5 series models pruned by depth pruning).
(a) Loss curves derived by P2 Law for validating generalization on dataset size (Qwen-2.5-0.5B pruned by width pruning).
(b) Loss curves derived by P2 Law for validating generalization on dataset size (Qwen-2.5-1.5B pruned by width pruning).
(c) Loss curves derived by P2 Law for validating generalization on dataset size (Qwen-2.5-3B pruned by width pruning).
@@
Figure 23: Loss curves derived by P2 Law for validating generalization on dataset size (Qwen-2.5 series models pruned by width pruning).
@@
(a) Loss curves derived by P2 Law for validating generalization on dataset size (Llama-3 series models pruned by 2:4 semi-structured pruning).
@@
(b) Loss curves derived by P2 Law for validating generalization on dataset size (Qwen-2.5 series models pruned by 2:4 semi-structured pruning).
@@
Figure 24: Loss curves derived by P2 Law of models pruned by 2:4 semi-structured pruning for validating generalization on dataset size.
(a) Comparison between loss curves predicted by P2 Law and actual loss curves for validating generalization on model size (Llama-3 series models pruned by depth pruning).
(b) Comparison between loss curves predicted by P2 Law and actual loss curves for validating generalization on model size (Qwen-2.5 series models pruned by depth pruning).
(c) Comparison between loss curves predicted by P2 Law and actual loss curves for validating generalization on model size (Qwen-2.5 series models pruned by width pruning).
Figure 25: Comparison between loss curves predicted by P2 Law and actual loss curves for validating generalization on model size.
(a) Comparison between loss curves predicted by P2 Law and actual loss curves for validating generalization on pruning rate (Llama-3 series models pruned by depth pruning).
(b) Comparison between loss curves predicted by P2 Law and actual loss curves for validating generalization on pruning rate (Llama-3 series models pruned by width pruning).
Figure 26: Comparison between loss curves predicted by P2 Law and actual loss curves for validating generalization on pruning rate (Llama-3 series models).
(a) Comparison between loss curves predicted by P2 Law and actual loss curves for validating generalization on pruning rate (Qwen-2.5 series models pruned by depth pruning).
(b) Comparison between loss curves predicted by P2 Law and actual loss curves for validating generalization on pruning rate (Qwen-2.5 series models pruned by width pruning).
Figure 27: Comparison between loss curves predicted by P2 Law and actual loss curves for validating generalization on pruning rate (Qwen-2.5 series models).

### v3

@@ 5.2 Scaling Law
We are the first to investigate the scaling law for the post-training after model pruning, and we propose the P2 Law as a scaling law for this process. Compared to the OpenAI scaling law, the Chinchilla scaling law demonstrates superior performance (detailed in Appendix E). Therefore, we adopt the Chinchilla scaling law as the foundational parameterization for our P2 Law.
@@ 6 Conclusion
Due to constraints in GPU resources, the experiments conducted in this paper are restricted to LLMs with fewer than 8B parameters. Given the substantial increase in experimental costs for larger-scale models—for instance, training a 70B LLM with 1B tokens on 4 A800 GPUs would require approximately 1,000 hours—we intend to expand our experiments to larger models as soon as sufficient computational resources become available. This will enable us to further validate the applicability of the P2 Law across a broader range of model parameter scales.
Acknowledgments
This work is supported by the National Key Research & Develop Plan (2023YFF0725100) and the National Natural Science Foundation of China (62322214, U23A20299, U24B20144, 62172424, 62276270).
@@ B.2 Width Pruning
In the post-training process, We train this 2:4 sparse model pruned by SparseGPT. Inspired by LoRS Hu et al. (2025), during the post-training process, we combine the updated weight $\Delta\tilde{W}^{t}$ from each training iterate $t$ with the mask $M$ to obtain the weight after update $\tilde{W}^{t}$, ensuring the model’s sparsity:
@@
Appendix C Batch Size and Learning Rate Settings
Previous research indicates that the relationship between batch size and the number of model parameters is very weak McCandlish et al. (2018). Furthermore, OpenAI Scaling Law also utilize the same batch size for models with varying parameter counts. As a result, we apply a consistent and commonly used batch size of 262k tokens across models of different scales. Regarding the learning rate, OpenAI suggests that the optimal learning rate follows a logarithmic relationship with the size of the model parameters Kaplan et al. (2020). Based on their provided formula, the optimal learning rate for 8B models is calculated to be 2e-3, while for 0.5B models, it is 1.8e-3, indicating a minimal difference. Furthermore, our experiments reveal that the optimal learning rate for post-training of models ranging from 0.5B to 8B is approximately 2e-5. Therefore, we adopt a uniform learning rate across models of different scales.
Appendix D Additional Actual Loss Curves
The additional post-training loss curves for models pruned by width pruning or for the Qwen-2.5 series models are provided in Figures 89, 10 and 12.
Appendix E Comparison with OpenAI Scaling Law
@@
Appendix F Parameter Values of Fitted Parameterizations
@@
Appendix G Additional Loss Curves Derived by P2 Law
The additional loss curves derived by P2 Law are shown in the Figure 131415 and 16.
Appendix H Patterns of the Llama-3 Series Models in Terms of Width
@@
Appendix I Additional Generalization Loss Curves
We present the additional dataset size generalization predicted loss curves in the Figure 18192021 and 22, model size generalization predicted loss curves in the Figure 23 and pruning rate generalization predicted loss curves in the Figure 24 and 25.
@@ Appendix I Additional Generalization Loss Curves
Llama-3 series $\mathcal{L}_{1}$ 0.02 5.94 0.14 -1.57 0.23 -1.08 0.29 0.05 5.86 -2.52 -1.68 0.08 -0.97 0.38 38.26 0.87 2.49 26.53 0.37 0.05
@@
(a) Loss curves fitted with the P2 Law using the first 80% of checkpoints; the remaining 20% are used for validation. (Llama-3.2-1B pruned by depth pruning)
(b) Loss curves fitted with the P2 Law using the first 80% of checkpoints; the remaining 20% are used for validation. (Llama-3.2-3B pruned by depth pruning)
(c) Loss curves fitted with the P2 Law using the first 80% of checkpoints; the remaining 20% are used for validation. (Llama-3.1-8B pruned by depth pruning)
@@
Figure 18: Generalization of the P2 Law for Llama-3 series models pruned by depth pruning on dataset size.
(a) Loss curves fitted with the P2 Law using the first 80% of checkpoints; the remaining 20% are used for validation. (Llama-3.2-1B pruned by width pruning)
(b) Loss curves fitted with the P2 Law using the first 80% of checkpoints; the remaining 20% are used for validation. (Llama-3.2-3B pruned by width pruning)
@@
Figure 19: Generalization of the P2 Law for Llama-3 series models pruned by width pruning on dataset size.
(a) Loss curves fitted with the P2 Law using the first 80% of checkpoints; the remaining 20% are used for validation. (Qwen-2.5-0.5B pruned by depth pruning)
(b) Loss curves fitted with the P2 Law using the first 80% of checkpoints; the remaining 20% are used for validation. (Qwen-2.5-1.5B pruned by depth pruning)
(c) Loss curves fitted with the P2 Law using the first 80% of checkpoints; the remaining 20% are used for validation. (Qwen-2.5-3B pruned by depth pruning)
@@
Figure 20: Generalization of the P2 Law for Qwen-2.5 series models pruned by depth pruning on dataset size.
(a) Loss curves fitted with the P2 Law using the first 80% of checkpoints; the remaining 20% are used for validation. (Qwen-2.5-0.5B pruned by width pruning)
(b) Loss curves fitted with the P2 Law using the first 80% of checkpoints; the remaining 20% are used for validation. (Qwen-2.5-1.5B pruned by width pruning)
(c) Loss curves fitted with the P2 Law using the first 80% of checkpoints; the remaining 20% are used for validation. (Qwen-2.5-3B pruned by width pruning)
@@
Figure 21: Generalization of the P2 Law for Qwen-2.5 series models pruned by width pruning on dataset size.
@@
(a) Loss curves fitted with the P2 Law using the first 80% of checkpoints; the remaining 20% are used for validation. (Llama-3 series models pruned by 2:4 semi-structured pruning)
@@
(b) Loss curves fitted with the P2 Law using the first 80% of checkpoints; the remaining 20% are used for validation. (Qwen-2.5 series models pruned by 2:4 semi-structured pruning)
@@
Figure 22: Generalization of the P2 Law for models pruned by 2:4 semi-structured pruning on dataset size.
(a) P2 Law is fitted using checkpoints from smaller LLMs and used to predict the loss curves of larger LLMs. (Llama-3 series models pruned by depth pruning)
(b) P2 Law is fitted using checkpoints from smaller LLMs and used to predict the loss curves of larger LLMs. (Qwen-2.5 series models pruned by depth pruning)
(c) P2 Law is fitted using checkpoints from smaller LLMs and used to predict the loss curves of larger LLMs. (Qwen-2.5 series models pruned by width pruning)
Figure 23: Generalization of the P2 Law on model size.
(a) P2 Law is fitted using checkpoints from smaller pruning rates and used to predict the loss curves of larger ones. (Llama-3 series models pruned by depth pruning)
(b) P2 Law is fitted using checkpoints from smaller pruning rates and used to predict the loss curves of larger ones. (Llama-3 series models pruned by width pruning)
Figure 24: Generalization of the P2 Law for Llama-3 series models on pruning rate.
(a) P2 Law is fitted using checkpoints from smaller pruning rates and used to predict the loss curves of larger ones. (Qwen-2.5 series models pruned by depth pruning)
(b) P2 Law is fitted using checkpoints from smaller pruning rates and used to predict the loss curves of larger ones. (Qwen-2.5 series models pruned by width pruning)
Figure 25: Generalization of the P2 Law for Qwen-2.5 series models on pruning rate.